OpenStudy (anonymous):

Our balance is maintained, at least in part, by the endolymph fluid in the inner ear. Spinning displaces this fluid, causing dizziness. Suppose a dancer (or skater) is spinning at a very high rate of 3.00 revolutions/second about a vertical axis through the center of his head. Although the distance varies from person to person, the inner ear is approximately 7.0 {\rm cm} from the axis of spin.What is the radial acceleration (in {\rm m}/{\rm s}^{2}) of the endolymph fluid?

OpenStudy (anonymous):

\[\omega = 2 \pi f\] \[\omega = 6 \pi\] \[a = \omega ^{2} R\] \[a = 36 \pi ^{2} R\] Convert R from rm/cm in rm/m and multiply it with 36\[\pi ^{2}\]

OpenStudy (anonymous):

but how u come to know we have totake max.acceleration?

OpenStudy (anonymous):

that's the best solution i have :p we got f = 3.00 rv/sec and for as long as f doesn't change that's the answer... because \[\pi\] and R are constants...